OpenStudy (anonymous):

Solve the equation using factoring x2 + 7x + 12 = 0 A. {2, 6} B. {-2, -6} C. {3, 4} D. {-3, -4}

hero (hero):

x^2+7x+12 = 0 x^2+3x+4x+12=0 x(x+3)+4(x+3)=0 (x+3)(x+4)=0 x = -3,-4

OpenStudy (saifoo.khan):

So D.

OpenStudy (nenadmatematika):

correct :D